# Pz Cats: Photometric redshift catalogs based on DES Y3 BAO sample

### v1

@@ Abstract
The photometric redshift estimation (photo-z) has been developed over the years with various methods. In this work, we analyse four different photo-z estimators using the Dark Energy Survey Y3 BAO Sample: ANNz2, BPZ, ENF, and DNF. Unlike what is usually found in the literature, we investigate the possibility of selecting the best galaxies according to their redshift Probability Distribution Function (PDF). We selected 25,760 galaxies from four different spectroscopic surveys and cross-matched them with the photo-z sample. These galaxies served to understand the redshift bias and its 68th percentile $\sigma_{68}$. We found that within a range of $0.850.79<z_{p}<$ 0.85 there is the lowest $\sigma$ for all the estimators we analysed. DNF has the biggest absolute value of the bias ($\sigma$), while ENF, ANNz2 and BPZ lose precision for a redshift range below 0.7 and higher than 0.9. If one wants to pick the best galaxies by removing the bins with the worst bias, one will find that ANNz2 is the most robust algorithm for all chosen criteria. When selecting the best PDFs, the resulting sub-samples gave BPZ with more selected objects. ANNz2 shows better precision, ENF has the worst selection of Gaussian PDFs, with very few galaxies left for an LSS study. We also showed that even though the PDFs are smooth, there are catastrophic redshift results. Lastly, DNF is the worst in precision but with sufficient galaxies for cosmological analysis. We also selected galaxies whose PDFs have only secondary peaks not bigger than 30% of the main peak height, called Small Peaks. For these sub-samples, ANNz2 outperformed the other algorithms. We will make all catalogs publicly available through the package Pz Cats.
@@ 1 Introduction
There are many thorough photometric redshift verifications made by the DES Collaboration or independent groups like [15, 16, 17]. Our main goal in this paper is to analyse how selecting galaxies through their full redshift Probability Distribution Function (PDF) can affect the effective redshift ($z_{eff}$) of the survey, its precision and sample size. This is different than assuming the PDFs are nearly Gaussian and comparing different photo-z estimators by their redshift variance. Furthermore, we are using a different training spectroscopic sample, with intersections with the ones used by the DES Collaboration, less populated, but still relevant to understand how each estimator can be influenced by the size of the training sample. In terms of precision, we obtain the photo-z errors without assuming any distribution function to the sample’s distribution $N(z)$.
@@ 2.2 ANNz2
It is possible to obtain photo-zs through an Artificial Neural Network (ANN), a training-based method. It forms a map between the input variables and output ones, the connections through response functions are called neurons. ANNz2 [18] is a software that uses a multilayer perceptron (MLP). Of the many layers of nodes, the inputs that are the magnitudes of $k$ filters, $\bf{m}_{k}$, are in the first layer, while the last ones contain the outputs: the photometric redshift $z_{p}$ and the PDFs.
To obtain $z_{p}$, the data must be divided into training, testing, and evaluation sets. For samples of galaxies without spectroscopic redshift, one needs to match the galaxies from a spectroscopic survey to use for training and testing.
@@ 2.4 Spectroscopic sample
In this work, we used the Dark Energy Survey BAO sample of red galaxies from the third year of observations, Dark Energy Survey (DES) Year 3 [19] photometric sample. The survey’s sample consists of $7,081,993$ galaxies. The input of each object was used in the three photo-z estimators within a range of $2.00<z<$ 2.0. We test four different algorithms for estimating the Probability Density Function: ANNz2, BPZ, ENF, DNF  [^1] We included the codes used here at: https://github.com/psilvaf/cat_org..
The reference spectroscopic galaxy set consists of $25,760$ galaxies from: the VIPERS Data Release 2 (PDR-2) [20] W1 and W4 equatorial fields, DEEP2 Galaxy Redshift Survey [21], VIMOS VLT Deep Survey (VVDS) [22], and the SDSS eBOSS LRG pCMASS [23]. These galaxies were found by matching their celestial coordinates with a tolerance of $2$ arcseconds. We used the Random Regression technique, dividing the reference set into four parts, the training with $20,608$ galaxies and three test sets each with $1,000$; $2,000$; and $2,760$ galaxies, respectively. An ensemble of $100$ random Machine Learning Methods was used in the photo-z estimation, $90$ kNN, and $100$ PDF bins. In Figure 1, we show the distribution of colours $g-i$ and $r-i$ for the DES observations for its whole BAO sample in black and our matched galaxies in blue, we do not have information for $r-i>1.5$, but this is shown to be irrelevant in the results. Figure 2 shows the redshift distribution of this sample, ideal for a LSS survey with LRG.
Figure 1: Colours $r-i$ and $g-i$ of the DES galaxies (black) and the matched spec-z sample (blue).
@@
In Table (1), we show the effective redshift $z_{eff}$ defined in [24] and the 68th percentile of the redshift error $\sigma_{z}$, defined as
@@
Table 1: Each survey with its effective redshift $z_{eff}$.
Figure 3: $\sigma_{68}$ of each photo-z estimator. The DNF’s bias is shown in the right y-axis.
@@ 3 Bias per redshift bin
We plot the bias, Eq. (3.1), per redshift bin in Figure (3), the sample used to compare is the same of the previous plots. The red squares represent the ENF bias per bin, they are larger for $z>0.95$, this is expected once we do not have enough galaxies with these redshift values. The same is true for ANNz2, but the bias is smaller indicating a better performance of this estimator. DNF is the most biased estimator for most bins. Compared to the DES Collaboration results in [15], we agree that in the bins around $z_{p}=0.8$, the bias increases at $z_{p}>1.0$.
In Table (1), we show the effective redshift $z_{eff}$ described by [24], and the 68th percentile of the redshift error $\sigma_{z}$, defined as
$\sigma_{z}=\frac{(z_{p}-z)}{(1+z)}.$ (3.2)
The comparison was made with the spectroscopic dataset used from training and testing. For ANNz2, we used the set separated to test for the algorithm pipeline. The BPZ bias was calculated with the available spectroscopic observations. Lastly, for ENF/DNF, we had to run the estimator with the galaxies of the training set, since ENF/DNF does not have a testing step. We separated 80% for training and 20% for testing ENF/DNF, the 20% was used to compute the bias between redshifts.
@@ 4 Results
Our results for the four methods are available at the Zenodo repository with DOI: 10.5281/zenodo.14290701. In Figure 4, we show the ANNz2 testing sample compared to the resulting photo-z. The cyan dotted line shows the $\sigma_{99}$ error, the dashed blue line represents the $\sigma_{68}$ error with the expected result. The black and white pixels represent the number of galaxies concentrated with that particular spec-z/photo-z value. The resulting redshift estimation was satisfactory, the sample distribution matches the spectroscopic redshift pattern without much dispersion.
Figure 4: ANNz2 photo-z compared to the spectroscopic training set. The horizontal axis is the spec-z testing sample.
Figure 5 shows the high bias between the BPZ photo-z and the matched samples $spec-z$. Compared to ANNz2, this estimator has a bigger error, indicated by the wider $\sigma_{68}$ region. The results are consistent with the expected redshift, with little dispersion from the $z_{p}=z_{spec}$ line. For redshifts higher than 1.1, $\sigma_{99}$ deviates from a straight line. The stripped pattern is due to the limitation of the estimator of setting a redshift resolution, we chose a resolution of $0.02$. This same pattern can be seen in other studies using BPZ, like in [25].
Figure 5: BPZ vs Vipers redshift results.
@@
In table 2, we included results using DES data with different algorithms, from the Scientific Verification and Deep Field. We had to estimate the $\sigma_{68}$ from [17] to the equivalent version of our definition. Our results show lower $\sigma_{68}$ than most cases. The exception is DNF Complete training (all the available filters from DECAM) set from [17]. Compared to [17]’s DNF incomplete case (only $g,r,i,z$ filters like the BAO sample), we found compatible results.
Figure 6: ENF photo-z vs Vipers redshift results.
Figure 7: DNF photo-z vs Vipers redshift results.
In order to make sure our spectroscopic sample was enough for DNF, we compare, in Figure 8, our result in the x-axis with the BAO sample by the DES Y3 [19] in the y-axis. They are highly correlated for all the redshift values. The DES Collaboration used $\sim 2.2\times 10^{5}$ spectroscopic matched galaxies in their estimation spectra matched [19]. Despite their superior access to spectroscopic data, our analysis is comparable with $\langle\sigma\rangle=z_{Pz\,\,cats}-z_{DESY3}=-0.003$. Figure 8 shows the 68th percentile for this bias with the same colour scheme used above. The results diverge for $z>0.9$, this means DNF is more sensitive to the training set in this range.
Figure 8: DNF results with our training set and the DESY3 one.
@@ 4.1 Consistency tests based on the redshift Probability Distribution Function
Photo-z estimator Method No. of galaxies $z_{eff}$ $\sigma_{z}^{68}$
ANNz2 Gaussian PDFs $2,931,677$ $0.817$ ${}^{+0.014}_{-0.031}$
ANNz2 Small peaks $5,133,775$ $0.799$ ${}^{+0.009}_{-0.022}$
ENF Gaussian PDFs $299,427$ $0.834$ ${}^{+0.040}_{-0.084}$
ENF Small peaks $2,256,616$ $0.768$ ${}^{+0.046}_{-0.055}$
DNF Gaussian PDFs $2,111,461$ 0.868 ${}^{+0.116}_{-0.121}$
DNF Small peaks $2,456,190$ $0.761$ ${}^{+0.033}_{-0.142}$
BPZ Gaussian PDFs $3,973,671$ $0.812$ ${}^{+0.016}_{-0.045}$
BPZ Small peaks $5,904,860$ $0.811$ ${}^{+0.008}_{-0.042}$
@@
The resulting sub-samples are shown in the table 3. BPZ has the biggest number of selected objects. However, because of the Bayesian method, this estimator forces smoother PDFs, but that may not mean it is the best representation of all galaxies, we need to trust that the template available is representative enough of the galaxy’s photometry. ANNz2 shows better precision despite not having a sample as big as BPZ’s. ENF has the worst selection of Gaussian PDFs, with very few galaxies left for a LSS study. Lastly, the worst in precision is DNF but with sufficient galaxies for cosmological analysis. The Small Peaks results are more populated, this time ANNz2 has better precision and more objects than the other algorithms. Choosing the least noisy PDFs may be an advantage to LSS studies for being less selective because the decrease in the number of objects is lower.
From eq. (3.1) 68th percentile, we plot the distribution of $\sigma_{z}$ for all the estimators and sample cuts. In Figure 9, we show the distribution of $\sigma_{z}$ for all the samples described here. The black line represents the whole survey, the blue one represents the samples only with Gaussian PDFs, and the grey area includes multi-modal PDFs but not too noisy.
@@
Figure 9: Comparison of $\sigma_{z}$ values from different models.
The first panel has the ANNz2 normalised bias, Figure 9(a), the vertical lines respecting the distribution colours. The expected normalised bias is a Gaussian distribution with mean zero, the three cases show the centres close to zero, but the distributions are not perfectly Gaussian, and the ±$\pm\sigma_{68}$ are not symmetric. The least accurate sample is the blue (Gaussian) one, the main reason is the loss of statistics. In Figure 9(b), BPZ’s three distributions are very similar and not Gaussian. ENF and DNF (Figure 9(c),9(d)) have the worst normalised bias distribution and biggest ±$\pm\sigma_{68}$.

### v3

@@ Abstract
Over the years, photometric redshift estimation (photo-z) has advanced through various methods. This study evaluates four distinct photo-z estimators—ANNz2, BPZ, ENF, and DNF—using the Dark Energy Survey Y3 BAO Sample. Unlike most studies, we explore selecting optimal galaxies based on their redshift Probability Distribution Function (PDF) by either reducing noise or identifying those approximating a Gaussian distribution. We cross-matched 25,760 galaxies drawn from four spectroscopic surveys with the photo-z sample to comprehend redshift bias and its 68th percentile $\sigma_{68}$. The lowest $\sigma$ for all estimators was found in the range $0.850.79<z_{p}<$ 0.85. Among the estimators, DNF exhibited the greatest bias, while ENF, ANNz2, and BPZ showed decreased precision outside 0.7 to 0.9 redshift range. To select galaxies with minimal bias, ANNz2 emerged as the most reliable algorithm across all criteria. PDFs selection significantly improves colour representation over the spectroscopic sample, underscoring the role of magnitude space in selection. While ANNz2 achieved superior precision, ENF poorly selected Gaussian PDFs, leaving few galaxies for LSS evaluation. Despite smooth PDFs, catastrophic redshift errors were present. Though DNF had the poorest precision, it offered enough galaxies for cosmological use. Subsampling galaxies with secondary peaks less than 30% of the main peak height, termed Small Peaks, showed ANNz2 excelled. The catalogs produced have been published as Pz Cats within the ZENODO repository.
@@ 1 Introduction
There are many thorough photometric redshift verifications made by the DES Collaboration or independent groups like [15, 16, 17]. Our main goal in this paper is to analyse how selecting galaxies through their full redshift Probability Distribution Function (PDF) can affect the effective redshift ($z_{\rm eff}$) of the survey, its precision and sample size. This is different than assuming the PDFs are nearly Gaussian and comparing different photo-z estimators by their redshift variance. Furthermore, we are using a different training spectroscopic sample, with intersections with the ones used by the DES Collaboration, less populated, but still relevant to understand how each estimator can be influenced by the size of the training sample. In terms of precision, we obtain the photo-z errors without assuming any distribution function to the sample’s distribution $N(z)$.
@@ 2 Photo-z estimators
In this work, we used the Dark Energy Survey BAO sample of red galaxies from the third year of observations, Dark Energy Survey (DES) Year 3 [18] photometric sample. The dataset collected from the survey comprises a total of $7,081,993$ galaxies, spread over an area of $21,078.79$ square degrees. This amounts to an observed density of approximately $333.60$ galaxies per square degree. The input of each object was used in the three photo-z estimators within a range of $2.00<z<$ 2.0. We test four different algorithms for estimating the Probability Density Function: ANNz2, BPZ, ENF, DNF  [^1] We included the codes used here at: https://github.com/psilvaf/cat_org..
@@ 2.2 ANNz2
It is possible to obtain photo-zs through an Artificial Neural Network (ANN), a training-based method. It forms a map between the input variables and output ones, the connections through response functions are called neurons. ANNz [19] is a software that uses a multilayer perceptron (MLP). Of the many layers of nodes, the inputs that are the magnitudes of $k$ filters, $\bf{m}_{k}$, are in the first layer, while the last ones contain the outputs: the photometric redshift $z_{p}$ and the PDFs.
To obtain $z_{p}$, the data must be divided into training, testing, and evaluation sets. For samples of galaxies without spectroscopic redshift, one needs to match the galaxies to a spectroscopic survey to use for training and testing.
@@ 2.4 Spectroscopic sample
The reference spectroscopic galaxy set consists of $25,760$ galaxies from: the VIMOS Public Extragalactic Redshift Survey (VIPERS) Data Release 2 (PDR-2) (redshift flag Flag2absent2$\geq$ 2≥ 2) [20] W1 and W4 equatorial fields, DEEP2 Galaxy Redshift Survey (Flag Q$>$3) [21], VIMOS VLT Deep Survey (VVDS) [22], and the Sloan Digital Sky Survey (SDSS) eBOSS LRG pCMASS [23] (HIZ_LRG flag [24]) . These galaxies were found by matching their celestial coordinates with a tolerance of $2$ arcseconds. We used the Random Regression technique, dividing the reference set into four parts, the training with $20,608$ galaxies and three test sets each with $1,000$; $2,000$; and $2,760$ galaxies, respectively. An ensemble of $100$ random Machine Learning Methods was used in the photo-z estimation, $90$ kNN, and $100$ PDF bins. In Figure 1, we show the distribution of colours $g-i$ and $r-i$ for the DES observations for its whole BAO sample in black and our matched galaxies in blue, we do not have information for $r-i>1.5$, but this is shown to be irrelevant in the results. Figure 2 shows the redshift distribution of this sample, ideal for an LSS survey with LRG.
Figure 1: Colours $r-i$ vs $g-i$. of the DES galaxies (black) and the matched spec-z sample (blue).
@@
In Table (1), we show the effective redshift $z_{\rm eff}$ defined in [25] and the 68th percentile of the redshift error $\sigma_{z}$, defined as
@@
Table 1: Uncertainty ($\sigma_{z}^{68}$) and effective redshift ($z_{\rm eff}$) of each algorithm using the full survey.
Figure 3: Redshift bias $\langle\sigma\rangle$ w.r.t. averaged redshift $\langle z_{bin}\rangle$ in different bins of each photo-z estimator. ANNz2 (blue line with star symbols), ENF(red line with dots), BPZ (gray line with swuares), and DNF (dotted line). The DNF’s bias is shown in the right y-axis. The error bars were estimated from the 68th percentile of $\langle\sigma\rangle$ distribution around zero.
@@ 3 Bias per redshift bin
We plot the bias, Eq. (3.1), per redshift bin in Figure (3) and their respective uncertainty based on the percentile 68 pra the bias distribution, the sample used to compare is the same as the previous plots. The error bars represent well the number of spectroscopic redshifts available per bin; they are smaller for bins with more spec-z. Consequently, the bias and errors are larger for higher redshift, a common lack of available data due to the nature of spectroscopic surveys. The red squares represent the ENF bias per bin, they are larger for $z>0.95$, this is expected once we do not have enough galaxies with these redshift values. The same is true for ANNz2, but the bias is smaller, indicating a better performance of this estimator. DNF is the most biased estimator for most bins. Compared to the DES Collaboration results in [15], we agree that in the bins around $z_{p}=0.8$, the bias increases at $z_{p}>1.0$.
In Table (1), we show the effective redshift $z_{\rm eff}$ described by [25], and the 68th percentile of the redshift error $\sigma_{z}$ 2.4. This first result is based on the full sample. The comparison was made with the spectroscopic data set used in the training and testing. For ANNz2, we used the separated set to test the algorithm pipeline. The BPZ bias was calculated with the available spectroscopic observations. Lastly, for ENF/DNF, we had to run the estimator with the galaxies of the training set, since ENF/DNF does not have a testing step. We separated 80% for training and 20% for testing ENF/DNF, the 20% was used to compute the bias between redshifts.
@@ 4 Results
Our results for the four methods are available in the Zenodo repository with DOI: 10.5281/zenodo.14290701. In Figure 4, we show the ANNz2 testing sample compared to the resulting photo-z. The cyan dotted line shows the $\sigma_{99}$ error, the dashed blue line represents the $\sigma_{68}$ error with the expected result. The black and white pixels represent the number of galaxies concentrated with that particular spec-z/photo-z value. The resulting redshift estimation was satisfactory, the sample distribution matches the spectroscopic redshift pattern without much dispersion.
Figure 4: ANNz2 photo-z compared to the spectroscopic training set. The horizontal axis is the spec-z testing sample. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
Figure 5 shows the high bias between the BPZ photo-z and the matched samples $spec-z$. Compared to ANNz2, this estimator has a bigger error, indicated by the wider $\sigma_{68}$ region. The results are consistent with the expected redshift, with little dispersion from the $z_{p}=z_{spec}$ line. For redshifts greater than 1.1, $\sigma_{99}$ deviates from a straight line. The stripped pattern is due to the limitation of the estimator of setting a redshift resolution, we chose a resolution of $0.02$. This same pattern can be seen in other studies using BPZ, like in [26].
Figure 5: BPZ vs spec-z full reference set. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
@@
In table 2, we included results using DES data with different algorithms, from the Scientific Verification and Deep Field. We had to estimate the $\sigma_{68}$ from [16] to the equivalent version of our definition  [^2] Considering their mean spec-z is around 0.75, we divide their $\sigma_{68}$ by (1+z).. Our results show lower $\sigma_{68}$ than most cases. The exception is DNF Complete training (all the available filters from DECAM) set from [17]. Compared to [17]’s DNF incomplete case (only $g,r,i,z$ filters like the BAO sample), we found compatible results.
Figure 6: ENF photo-z vs spec-z. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
Figure 7: DNF photo-z vs spec-z. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between photo-z and spec-z. The blue dashed-line represents the 68% confidence level in a $\sigma_{z}^{68}$ distribution, and the cyan-dotted one the 99% confidence level.
In order to make sure our spectroscopic sample was enough for DNF, we compare, in Figure 8, our result in the x-axis with the BAO sample by the DES Y3 [18] in the y-axis. They are highly correlated for all the redshift values. The DES Collaboration used $\sim 2.2\times 10^{5}$ spectroscopic matched galaxies in their estimation spectra matched [18]. Despite their superior access to spectroscopic data, our analysis is comparable with $\langle\sigma\rangle=\langle z_{Pz\,\,cats}-z_{DESY3}\rangle=-0.003$. Figure 8 shows the 68th percentile for this bias with the same colour scheme used above. The results diverge for $z>0.9$, this means DNF is more sensitive to the training set in this range.
Figure 8: DNF results with our training set vs the DESY3 one. The colorbar represents the number of galaxies in a 2D histogram. The cyan solid-line is the expected linear relation between the two data sets. The blue dashed-line represents the 68% confidence level in a $z_{DES}-z_{PzCats}$ distribution, and the cyan-dotted one the 99% confidence level.
@@ 4.1 Consistency tests based on the redshift Probability Distribution Function
Estimator Method No. of galaxies $z_{\rm eff}$ $\sigma_{z}^{68}$ Area (deg2) Galaxies/deg2
ANNz2 Gaussian $2,931,677$ $0.817$ ${}^{+0.014}_{-0.031}$ 21,065.36 139.17
ANNz2 Small peaks $5,133,775$ $0.799$ ${}^{+0.009}_{-0.022}$ 21,065.36 243.71
ENF Gaussian $299,427$ $0.834$ ${}^{+0.040}_{-0.084}$ 21,032.17 14.24
ENF Small peaks $2,256,616$ $0.768$ ${}^{+0.046}_{-0.055}$ 21,065.36 107.12
DNF Gaussian $2,111,461$ 0.868 ${}^{+0.116}_{-0.121}$ 21,078.79 100.17
DNF Small peaks $2,456,190$ $0.761$ ${}^{+0.033}_{-0.142}$ 21,078.79 116.52
BPZ Gaussian $3,973,671$ $0.812$ ${}^{+0.016}_{-0.045}$ 21,065.36 188.64
BPZ Small peaks $5,904,860$ $0.811$ ${}^{+0.008}_{-0.042}$ 21,065.36 280.31
@@
Contrasting with the reweighing approach outlined in [28, 29], which constructs the probability density function (PDF) of the sample by using the weights while also avoiding both the need for trimming and the evaluation of the PDFs for individual galaxies, our sample selection methodology is specifically structured to remove unreliable individual PDFs that might exist within the survey dataset. These PDFs are prone to noise due to their under-representation in the magnitude space, leading to potential distortions in data interpretation.
The resulting sub-samples are shown in the table 3. BPZ has the biggest number of selected objects in both criteria. However, because of the Bayesian method, this estimator forces smoother PDFs, but that may not mean it is the best representation of all galaxies, we need to trust that the template available is representative enough of the galaxy’s photometry. ANNz2 shows better precision despite not having a sample as big as BPZ’s. ENF has the worst selection of Gaussian PDFs, with very few galaxies left for an LSS study. Lastly, the worst in precision is DNF but with sufficient galaxies for cosmological analysis. The Small Peaks results are more populated, this time ANNz2 has better precision and more objects than the other algorithms. Choosing the least noisy PDFs may be an advantage to LSS studies for being less selective because the decrease in the number of objects is lower.
The next question is to determine if the selected samples continue to represent robust large-scale structure (LSS) datasets. In Table 3, we have presented both the residual survey area and the density of objects, expressed as the number of galaxies per square degree. It is important to highlight that the spatial footprint of the survey has remained consistent, unaffected by the decrease in the quantity of objects. Nonetheless, there is a crucial caveat regarding the ENF Gaussian sample: the substantial reduction in object count for this particular sample renders it unsuitable for further analyses involving angular power spectra or two-point correlation functions. Our analysis of the galactic distribution through various cuts supports the data in Table 3. The ENF approach shows a decrease in galaxy counts, while an optimal sky direction features increased galaxy density linked to the spec-z sample. In contrast, ANNz2 exhibits the smallest sky distribution post-cuts.
The principal aim of this survey is to identify the BAO signal. However, a reduction in the number of galaxies compromises the signal, and the of this reduction on statistical power in [30]. It is important to highlight that BAO detection can be accomplished using two methodologies: one that operates in three dimensions and another that is transverse to the line of sight. To achieve accurate transverse BAO measurements, it is crucial to have an adequate density of galaxies per square degree. In contrast, for three-dimensional observations, the redshift distribution needs to be robust enough to capture the characteristic feature along the line of sight, irrespective of whether the survey footprint encompasses an extensively filled section of the sky. In particular, the most recent findings from the Dark Energy Spectroscopic Instrument (DESI) [31] provide an exemplary estimation of 3D BAO even when significant gaps are present across all its tracers. In the last column of Table 3, we included the number of galaxies per deg2, the only sample to lose statistics considerably is ENF, all the others are capable of being a BAO sample.
(a) BPZ
(b) ANNz2
(c) ENF
(d) DNF
Figure 9: Comparison of colour distributions for different methods of photo-z (black) estimation and the matched spec-z sample (blue). The contours represent the $68\%$ confidence level of the distribution: (a) BPZ, (b) ANNz2, (c) ENF, and (d) DNF.
(a) BPZ
(b) ANNz2
(c) ENF
(d) DNF
Figure 10: Comparison of colour distributions for different methods of photo-z (black) estimation and the matched spec-z sample (blue). The contours represent the $68\%$ confidence level of the distribution: (a) BPZ, (b) ANNz2, (c) ENF, and (d) DNF.
[…]
From eq. (3.1) 68th percentile, we plot the distribution of $\sigma_{z}$ for all estimators and sample cuts. In Figure 12, we show the distribution of $\sigma_{z}$ for all the samples described here. The black line represents the whole survey, the blue one represents the samples only with Gaussian PDFs, and the grey area includes multimodal PDFs but not too noisy.
@@
Figure 12: Comparison of $\sigma_{z}$ values from different models/algorithms. The vertical red-line marks the null difference, the black histogram has all galaxies, gray histogram represents the Small peaks cuts, and the blue histogram the Gaussian one with the corresponding colours of dashed vertical lines representing the 68% confidence level.
The first panel has the ANNz2 normalised bias, Figure 12(a), the vertical lines respecting the distribution colours. The expected normalised bias is a Gaussian distribution with mean zero, the three cases show the centres close to zero, but the distributions are not perfectly Gaussian, and the ±$\pm\sigma_{68}$ are not symmetric. The least accurate sample is the blue (Gaussian) one, the main reason is the loss of statistics. In Figure 12(b), BPZ’s three distributions are very similar and not Gaussian. ENF and DNF (Figure 12(c),12(d)) have the worst normalised bias distribution and biggest ±$\pm\sigma_{68}$.